4.48L of an ideal gas at STP requires 12 cal to raise the temperature by 15° C at constant volume. Find the Cp of the gas.

Grade:12

1 Answers

APURV GOEL
12 years ago

We have nCvΔT = ΔH

              4.48/22.4 x Cv x 15 = 12

     so Cv = 4

and Cp = Cv + r

           = 4 + 2 = 6

If u like my ans, please approve it.